OpenStudy (anonymous):

A car is driving at a speed of 60 mi/h. What is the speed of the car in feet per minute?

OpenStudy (ipwnbunnies):

We need to use conversion factors in this problem. These are the ones you need: \[\frac{5280 feet}{1 mile}\] \[\frac{1 hour}{60 minutes}\]

OpenStudy (ipwnbunnies):

There are 5280 feet in a mile and there are 60 minutes in an hour. We want the units to cancel out so we will be left with feet/min.

OpenStudy (johnweldon1993):

\[\large \frac{60 \text{Miles}}{1\text{Hour}} \times \frac{1 \text{Hour}}{60 \text{Minutes}} \times \frac{5280 \text{Feet}}{1 \text{Mile}}\] So all that equals \[\large \frac{1 \text{Mile}}{1 \text{Minute}} \times \frac{5280 \text{Feet}}{1\text{Mile}}\] And then we get down to \[\large \frac{5280 \text{Feet}}{\text{Minute}}\]